A U.S. firm sells merchandise today to a Japanese company for ¥100,000,000. The current exchange rate is ¥110.58/$, the account is receivable in three months, and the firm chooses to avoid any hedging techniques designed to reduce or eliminate the risk of changes in the exchange rate. If the exchange rate changes to ¥110.13/$, the U.S. firm will realize a ________ of ________.

a) translation gain; ¥3,695

b) transaction gain; $3,695

c) None of the above is correct
Answer #1

1 $ = Yen 110.58

OR 1 Yen = 1/ 110.58 = $ 0.00904322662

Forward Rate : 1 $ = Yen 110.13

OR 1 Yen = 1/ 110.13 = $ 0.00908017797

Transalation Gain** = (0.00908017797 - 0.00904322662) * 100,000,000

= 0.00003695135 * 100,000,000

= $ 3695

** This results in transaction Gain because there flow is involved. If there were only change in value of Asset or liability, it would have resulted in translation Gain.

Answer : - Transaction Gain of  $ 3695

Option B is correct.